Show that a function w(x, y) = cos(2x + 2ct) satisfies wave equation.
Answer #1

= cos (9x+2ct) W(xiy ) derivative Iw wirot it - sin (22+2(t) Fot = -sin (22+act) · (0+2c) 26 at (22+Qct) ow 3-2 sin(ay+act)